OpenStudy (anonymous):

What is the distance between the points (-4, 1) and (6, 8)? Round your answer to the nearest hundredth.

OpenStudy (anonymous):

pythagoras here

OpenStudy (anonymous):

from -4 to 6 is 10 from 1 to 8 is 7 so answer is \[\sqrt{10^2+7^2}=\sqrt{149}\]

OpenStudy (anonymous):

hey what was thfacetime geek talking about

OpenStudy (anonymous):

@estudier nice. hopefully you have started a trend and a) people will learn how to do this for themselves to make sure they get the right answer b) people here will stop showing off by posting solutions they find on wolfram,as if they can compute integrals requiring integration by parts twice in their heads in 10 seconds

OpenStudy (anonymous):

i mean if you just want a quick answer and no method, what better way?

OpenStudy (anonymous):

Yeah. Such a good point, Satellite. I know I've copied estudiar once the alpha link once or twice. And Satellite, thank you for pointing out that he should know and use the Pythagorean theorem instead of trying to teach him the "distance formula." The distance formula is so dumb.

OpenStudy (anonymous):

thank you, i will take that as a compliment

OpenStudy (anonymous):

Haha good. It was meant to be.